### v1

# Deterministic implementation in single-item auctions

Abstract
Deterministic auctions are attractive in practice due to their transparency, simplicity, and ease of implementation, motivating a sharp understanding of when they match randomized designs. We study deterministic implementation in single-item auctions under two outcome notions: (revenue, welfare) pairs and interim allocations. For (revenue, welfare) pairs, we show a discrete separation: there exists a pair implementable by a deterministic Bayesian incentive-compatible (BIC) auction but not by any deterministic dominant-strategy incentive-compatible (DSIC) auction. For continuous atomless priors, we identify conditions under which deterministic DSIC auctions are implementationally equivalent to randomized BIC auctions in terms of achievable outcomes. For interim allocations, we establish a deterministic analogue of Border’s theorem for two bidders, providing a necessary and sufficient condition for deterministic DSIC implementability, and use it to exhibit an interim allocation implementable by a deterministic BIC auction but not by any deterministic DSIC auction.
1 Introduction
The central objective of mechanism design is to identify and characterize the set of outcomes that can be implemented by suitable mechanisms. Since the foundational contributions of hurwicz1960optimality; hurwicz1972informationally, a substantial literature has developed across diverse market environments and under multiple solution concepts (clarke1971multipart; dobzinski2022hardness; feldman2023simultaneous; groves1973incentives; myerson1979incentive; myerson1981optimal; myerson1982optimal; vickrey1961counterspeculation; you2021incentive).
This paper advances that agenda in the setting of single-item auctions by studying the implementability of deterministic auctions, where allocation and payment rules are deterministic. Determinism is important in many applications, including antique auctions, spectrum sales, and platform marketplaces, where transparency, compliance, or operational simplicity may rule out randomization (klemperer2004auctions; liu2024optimal). Recent work on single-item auctions shows that, for revenue maximization, optimal auctions can often be chosen deterministic and DSIC in discrete settings, despite allowing randomized BIC mechanisms (giannakopoulos2024discrete). These optimality results, however, leave open a complementary question that is central to implementability: beyond the outcomes induced by optimizing a fixed objective, when does determinism preserve (or fail to preserve) the full set of achievable trade-offs and reduced forms? A precise account of the power and limitations of deterministic auctions is therefore both practically relevant and theoretically significant.
We take two complementary views. First, an outcome-space view that charts the set of achievable (revenue, welfare) pairs, together with its Pareto frontier. Second, a reduced-form view that characterizes the interim allocations implementable by deterministic mechanisms. These two views allow us to assess determinism at the level of aggregate trade-offs and at the level of allocation structure.
1.1 Implementation with respect to (revenue, welfare) pairs
Our first objective is to characterize which (revenue, welfare) pairs are attainable by deterministic DSIC auctions. For randomized auctions, the picture is well understood: one first identifies the extreme points and then takes convex combinations to generate the full attainable set. myerson1983efficient characterize the extreme points for the single-buyer case. likhodedov2003mechanism extend this to multiple buyers. This program does not carry over to deterministic auctions, because a convex combination of two deterministic auctions is not necessarily deterministic. As a result, the results and techniques of likhodedov2003mechanism do not apply, and the set of extreme points for deterministic auctions is not continuous.
We show that deterministic DSIC and deterministic BIC are not equivalent with respect to (revenue, welfare) pairs. There exists a pair that is implementable by a deterministic BIC auction but not by any deterministic DSIC auction. Prior results on revenue inequivalence between DSIC and BIC do not apply here, since they are confined to two-item environments (tang2016optimal; yao2017dominant). To further delineate the boundary of this separation, we identify a restricted class of distributions under which randomized BIC auctions and deterministic DSIC auctions remain equivalent with respect to (revenue, welfare) pairs. In the continuous case, we obtain two levels of equivalence: under regularity we match the Pareto frontier, and under stronger distributional conditions (in the i.i.d. case) we match the entire attainable region.
1.2 Implementation with respect to interim allocation
Our second objective is to characterize which interim allocations can be implemented by deterministic auctions. Here, “interim allocation” refers to the expected allocation probability, taken over the randomness in the other players’ types, in the same sense as defined in myerson1981optimal. border1991implementation gives a necessary and sufficient condition, now called Border’s theorem, for when an interim allocation (the “reduced form” in Border’s terminology) is implementable by a randomized BIC auction. Building on this, manelli2010bayesian show that any interim allocation of a randomized BIC auction is also implementable by a randomized DSIC auction. Further, chen2019equivalence prove that for continuous atomless priors, every interim allocation implementable by a randomized BIC auction is also implementable by a deterministic BIC auction. To the best of our knowledge, no analogous result is known for deterministic DSIC auctions. We show that the deterministic analogue fails: the equivalence does not extend to deterministic DSIC auctions. To show this separation, we introduce new techniques, most notably a deterministic version of Border’s theorem.
1.3 Our results
We study the power and limitations of deterministic auctions in single-item environments through two outcome notions: (revenue, welfare) pairs and interim allocation.
  • Revenue-welfare outcomes. We characterize which (revenue, welfare) pairs are attainable by deterministic DSIC auctions. In discrete settings, we exhibit a counterexample (via computational enumeration) showing a pair that is implementable by a deterministic BIC auction but by no deterministic DSIC auction, establishing a non-equivalence between the two. In contrast, under continuous regular priors, every Pareto-optimal (revenue, welfare) pair achievable by randomized BIC auctions is implementable by a deterministic DSIC auction. Moreover, in the i.i.d. continuous case, under Condition (1), every (revenue, welfare) pair implementable by randomized BIC auctions is also implementable by a deterministic DSIC auction.
  • Interim allocations. We provide a deterministic analog of Border’s theorem for two bidders, yielding necessary and sufficient conditions for when an interim allocation is implementable by a deterministic DSIC auction. Using this characterization, we show that deterministic BIC and deterministic DSIC are not equivalent. For three bidders with constant interim allocations, we further derive a necessary and sufficient condition for implementability under deterministic DSIC.
1.4 Related works
diakonikolas2012efficiency define the subset of (revenue, welfare) pairs that are undominated in both coordinates as Pareto points, and call their collection the Pareto curve, i.e., the upper-right boundary of the attainable set. They show that even with two buyers facing asymmetric discrete priors, this curve need not be concave, and that deciding whether a given point is achievable by a deterministic auction is NP-hard. In contrast, we prove that under continuous and regular priors, the Pareto frontier of randomized BIC mechanisms coincides with that of deterministic DSIC mechanisms.
arigapudi2018equivalence studies the related implementability problem in a symmetric environment. They give necessary and sufficient conditions under which the interim allocation of a symmetric BIC auction is implementable by a symmetric DSIC auction. We consider the two-buyer case without symmetry and our results specialize to theirs. Moreover, our theorem (Theorem 4) does not assume that the target interim allocation is implementable by any deterministic BIC mechanism. On the randomized side, border2007reduced extend Border’s theorem to asymmetric settings, and che2013generalized provide a clean network-flow proof. For deterministic implementability, bikhchandani2006weak show that a deterministic social choice function over multi-dimensional types is DSIC if and only if it is weakly monotone. gershkov2013equivalence further broaden equivalence results for randomized BIC and randomized DSIC mechanisms.
The study of attainable (revenue, welfare) outcomes is a recurring theme in auction theory (e.g., (bergemann2015limits; ma2022revenue)). hartline2008optimal design an auction that maximizes buyer surplus (welfare minus revenue), which corresponds to an extreme point of the outcome space. pai2009competing analyzes competition among sellers, yielding a revenue-utility trade-off for buyers. kleinberg2013ratio establish lower bounds on the revenue-to-welfare ratio under various feasibility constraints. In a similar spirit, bachrach2014optimising optimize a composite objective $\lambda_{1}$revenue + $\lambda_{2}$welfare + $\lambda_{3}$ click yield” to balance stakeholder goals, and shen2017practical propose a parametric family that optimizes a linear combination of revenue and welfare inside monotone transformations, achieving near-optimal revenue with practical implementability.
giannakopoulos2024discrete study discrete single-item optimal auction design and show that, for revenue maximization (and more generally linear objectives), an optimal mechanism can be chosen deterministic DSIC via an LP-duality/KKT characterization. Our paper is complementary: instead of identifying an optimizer for a fixed objective, we characterize which (revenue, welfare) trade-offs and interim allocations are implementable under determinism, and we obtain separations between deterministic BIC and deterministic DSIC. Since their techniques are tailored to optimality for a given objective, they do not imply our implementability questions or preclude the separations we establish.
2 Preliminaries
There are $n$ buyers competing for a single item. We use $n]=\{1,2,\cdots,n\}$ to represent the set of $n$ buyers. Each buyer $i$ has a private value $v_{i}\in\mathbb{R}_{\geq 0}$, drawn from a distribution $F_{i}$ with density function $f_{i}$. Valuations are independently drawn across buyers. We use $\boldsymbol{v}=(v_{1},v_{2},\cdots,v_{n})$ to denote the valuation profile of all $n$ buyers and $\boldsymbol{v}_{-i}=(v_{1},\cdots,v_{i-1},v_{i+1},\cdots,v_{n})$ to denote the valuation profile of all buyers except buyer $i$. Each buyer $i$ submits a bid $b_{i}\in\mathbb{R}_{\geq 0}$, and the bid profile is given by $\boldsymbol{b}=(b_{1},b_{2},\cdots,b_{n})$, with $\boldsymbol{b}_{-i}=(b_{1},\cdots,b_{i-1},b_{i+1},\cdots,b_{n})$ representing the bid profile of all buyers except buyer $i$. Throughout, we restrict attention to truthful (incentive-compatible) mechanisms. We use $x_{i}(\boldsymbol{v}):\mathbb{R}_{\geq 0}^{n}\rightarrow[0,1]$, $p_{i}(\boldsymbol{v}):\mathbb{R}_{\geq 0}^{n}\rightarrow\mathbb{R}_{\geq 0}$ to denote the ex-post allocation and the ex-post payment, respectively. An auction is deterministic if $x_{i}(\boldsymbol{v})\in\{0,1\}$ for any $i$ and any $\boldsymbol{v}$. With a slight abuse of notation, we use $x_{i}(v_{i})$, $p_{i}(v_{i})$ to denote interim allocation and interim payment, respectively. The intended meaning will be clear from context. $x_{i}(v_{i})$ and $p_{i}(v_{i})$ are defined as expectations over $\boldsymbol{v}_{-i}$. Specifically, the interim allocation can be expressed as $x_{i}(v_{i})=\int_{\boldsymbol{v}_{-i}}x_{i}(v_{i},\boldsymbol{v}_{-i})f_{-i}(\boldsymbol{v}_{-i})\mathrm{d}\boldsymbol{v}_{-i}$ where $f_{-i}(\boldsymbol{v}_{-i})=\prod_{j\neq i}f_{j}(v_{j})$ and $\mathrm{d}\boldsymbol{v}_{-i}=\mathrm{d}v_{1}\cdots\mathrm{d}v_{i-1}\mathrm{d}v_{i+1}\cdots\mathrm{d}v_{n}$. The interim payment can be expressed as $p_{i}(v_{i})=\int_{\boldsymbol{v}_{-i}}p_{i}(v_{i},\boldsymbol{v}_{-i})f_{-i}(\boldsymbol{v}_{-i})\mathrm{d}\boldsymbol{v}_{-i}$.
We consider the BIC, DSIC, and interim Individually Rational (IR) constraints in mechanisms design. A mechanism is BIC if truth-telling is the optimal strategy for each buyer, given that they have probabilistic beliefs about the valuations of other buyers. A mechanism is DSIC if truth-telling always maximizes a buyer’s utility, regardless of the actions of other buyers. A mechanism is interim IR if each buyer receives a non-negative expected utility from participating. Formally, these constraints are defined as follows:
  • BIC constraint: $v_{i}\cdot x_{i}(v_{i})-p_{i}(v_{i})\geq v_{i}\cdot x_{i}(b_{i})-p_{i}(b_{i})$, for all $i,v_{i},b_{i}\in\mathbb{R}_{\geq 0}$;
  • DSIC constraint: $v_{i}\cdot x_{i}(v_{i},\boldsymbol{b}_{-i})-p_{i}(v_{i},\boldsymbol{b}_{-i})\geq v_{i}\cdot x_{i}(b_{i},\boldsymbol{b}_{-i})-p_{i}(b_{i},\boldsymbol{b}_{-i})$, for all $i,v_{i},\boldsymbol{b}_{-i},b_{i}\in\mathbb{R}_{\geq 0}$;
  • Interim IR constraint: $v_{i}\cdot x_{i}(v_{i})-p_{i}(v_{i})\geq 0$, for all $i,v_{i}$.
According to myerson1981optimal, a BIC auction is characterized by the following conditions: $x_{i}(v_{i})$ is nondecreasing (monotone) in $v_{i}$ for all $i$, and the payment function satisfies the identity $p_{i}(v_{i})=v_{i}\cdot x_{i}(v_{i})-\int_{0}^{v_{i}}x_{i}(t)\mathrm{d}t+p_{i}(0)$. Under interim IR, we normalize $p_{i}(0)=0$.
According to nisan2007algorithmic, a DSIC auction is characterized by the following conditions: $x_{i}(v_{i},\boldsymbol{b}_{-i})$ is nondecreasing (monotone) in $v_{i}$ for all $i$ and any given $\boldsymbol{b}_{-i}$, and the payment function satisfies $v_{i}\cdot x_{i}(v_{i},\boldsymbol{b}_{-i})-\int_{0}^{v_{i}}x_{i}(t,\boldsymbol{b}_{-i})\mathrm{d}t$. Notably, any DSIC auction is also BIC.
The formulas for expected welfare and expected revenue are as follows:
$\textrm{WEL}=\sum_{i=1}^{n}\int_{v_{i}}v_{i}x_{i}(v_{i})f_{i}(v_{i})\mathrm{d}v_{i},\textrm{REV}=\sum_{i=1}^{n}\int_{v_{i}}\phi_{i}(v_{i})x_{i}(v_{i})f_{i}(v_{i})\mathrm{d}v_{i},$
where $\phi_{i}(v_{i})=v_{i}-\frac{1-F_{i}(v_{i})}{f_{i}(v_{i})}$, known as the “virtual value” in myerson1981optimal. We also use the quantile space (bulow1989simple). Let $q_{i}=F_{i}(v_{i})\in[0,1]$, and with a slight abuse of notation, let $v_{i}(q_{i})=F_{i}^{-1}(q_{i})$, the inverse function of $F_{i}(v_{i})$. Then, in the quantile space, we have
$\textrm{WEL}=\sum_{i=1}^{n}\int_{q_{i}}v_{i}(q_{i})\hat{x}_{i}(q_{i})\mathrm{d}q_{i},\textrm{REV}=\sum_{i=1}^{n}\int_{q_{i}}\hat{\phi}_{i}(q_{i})\hat{x}_{i}(q_{i})\mathrm{d}q_{i},$
where $\hat{\phi}_{i}(q_{i})=v_{i}(q_{i})-(1-q_{i})v_{i}^{\prime}(q_{i})$, and $\hat{x}_{i}(q_{i}):[0,1]\rightarrow[0,1]$ is the interim allocation when $F_{i}(v_{i})=q_{i}$ in the quantile space.
A point $P$ is a Pareto point of the range of the (revenue, welfare) pairs if there is no point $P^{\prime}\neq P$ in the range that dominates $P$ (coordinate-wise).
Implementability: a (revenue, welfare) pair is implemented by mechanism $\mathcal{M}$ if and only if $\mathcal{M}$ yields revenue exactly and welfare exactly; an interim allocation rule $\{x_{i}(\cdot)\}_{i\in[n]}$ is implemented by mechanism $\mathcal{M}$ if and only if $\mathcal{M}$ induces interim allocation exactly $\{x_{i}(\cdot)\}_{i\in[n]}$.
We present a variant of Border’s theorem (border1991implementation; che2013generalized) that, in our setting, characterizes exactly when an interim allocation is implementable by a randomized BIC auction.
Border’s theorem: an interim allocation profile $(x_{1}(\cdot),\cdots,x_{n}(\cdot))$ can be implemented by a randomized BIC auction if and only if
$\forall(v_{1},\cdots,v_{n}):~\sum_{i}\int_{t_{i}=v_{i}}^{\infty}x_{i}(t_{i})f_{i}(t_{i})\mathrm{d}{t_{i}}\leq 1-\prod_{i}F_{i}(v_{i}).$
Especially for a symmetric auction in the quantile space, we have
$\forall q_{i}:~\int_{t_{i}=q_{i}}^{1}\hat{x}_{i}(t_{i})\mathrm{d}t_{i}\leq\frac{1-q_{i}^{n}}{n}.$
3 Implementation with respect to (revenue, welfare) pairs
3.1 Nonequivalence of deterministic BIC auctions and deterministic DSIC auctions with respect to Pareto points in discrete case
In this subsection, we give an example in the i.i.d. discrete case to derive the following theorem.
Theorem 1.
There exists a Pareto point $P$ that is implemented by a deterministic BIC auction but cannot be implemented by any deterministic DSIC auction in i.i.d discrete case.
Suppose that there are two identical buyers, each with three types: $t_{1}=100,t_{2}=10,t_{3}=1$, occurring with probabilities $p_{1}=0.8,p_{2}=0.15,p_{3}=0.05$ respectively. For each type of buyer 1 and buyer 2, the item can be allocated to no one, to buyer 1, or to buyer 2. Thus, we have to check $3^{9}=19683$ deterministic allocation profiles. Similar to diakonikolas2012efficiency, to consider Pareto-optimal points, we can assume without loss of generality that payments are supported on the prior distribution. Therefore, we can enumerate all deterministic allocation rules on the finite type space and, for each, compute supporting payments and check BIC/DSIC/IR  [^1] Otherwise, there would be infinitely many feasible actions.. Through experiments, we find that $(\textrm{WEL}=96.2275,\textrm{REV}=85)$ is a Pareto-optimal point among all deterministically implementable outcomes (see Figure 1 in Appendix A.1).
The Pareto-optimal point $(\textrm{WEL}=96.2275,\textrm{REV}=85)$ can be achieved only by four deterministic BIC auctions, with $i\in\{1,2\}$ and $j\in\{1,2\}$, as shown in Table 1. However, none of these auctions is DSIC.
Table 1: Four deterministic BIC auctions implement the Pareto-optimal point $(\textrm{WEL}=96.2275,\textrm{REV}=85)$. Here the identification $v_{1}(v_{2})$ represents the type of buyer 1(2). The elements represent the buyer who gets the item, where “seller” means that no buyer gets the item.
$v_{2}$ $v_{1}$ $t_{3}=1$ $t_{2}=10$ $t_{1}=100$
$t_{3}=1$ buyer $j$ seller buyer 1
$t_{2}=10$ seller buyer $j$ buyer 1
$t_{1}=100$ buyer 2 buyer 2 buyer $i$
Note that in Table 1, buyer $j$ and buyer $i$ denote any buyer. Therefore, the four deterministic auctions are: (1) $i=1,j=1$; (2) $i=1,j=2$; (3) $i=2,j=1$; (4) $i=2,j=2$. Here, we use the case $i=2,j=1$ to illustrate that this deterministic auction is BIC but not DSIC. For buyer 1, we can get the interim allocation, that is $x_{1}(1)=0.05$, $x_{1}(10)=0.15$ and $x_{1}(100)=0.2$. For buyer 2, we can get the interim allocation, that is $x_{2}(1)=0$, $x_{2}(10)=0$ and $x_{2}(100)=1$. Clearly, the allocation functions of both buyers are monotonic. The welfare generated by buyer 1 is $0.05\times 1\times 0.05+0.15\times 10\times 0.15+0.2\times 100\times 0.8=16.2275$; the welfare generated by buyer 2 is $0\times 1\times 0.05+0\times 10\times 0.15+1\times 100\times 0.8=80$; the total welfare is $16.2275+80=96.2275$. As for revenue, we have to design payment rules for both buyers. However, unlike the case with continuous types, when types are discrete, there can be multiple possible payment rules that satisfy BIC. Based on the established allocation rule, we can formulate an optimization problem that maximizes the seller’s expected revenue while incorporating the BIC constraints as part of the feasible set. By solving this constrained optimization problem, we can derive the corresponding payment functions for both buyers. Finally, we can get the interim payment, that is $p_{1}(1)=0.05$, $p_{1}(10)=1.05$, $p_{1}(100)=6.05$, $p_{2}(1)=0$, $p_{2}(10)=0$, $p_{2}(100)=100$. The revenue generated by buyer 1 is $0.05\times 0.05+1.05\times 0.15+6.05\times 0.8=5$; the revenue generated by buyer 2 is $0\times 0.05+0\times 0.15+100\times 0.8=80$; the total revenue is $5+80=85$.
The allocation function clearly fails to satisfy the monotonicity requirement for DSIC, as the assignment does not increase monotonically with buyer types. For example, when buyer 1’s type is 1, as buyer 2’s type increases, the allocation transitions from buyer 1 to the seller and then to buyer 2. This transition pattern violates the monotonicity condition required for DSIC implementability. Hence, none of these four auctions can satisfy the dominant-strategy incentive-compatibility constraints, even though they all achieve the same Pareto-optimal (revenue, welfare) pair under BIC.
3.2 Equivalence of Deterministic DSIC and Randomized BIC Auctions with respect to Pareto Points For Regular Distribution
In this subsection, we assume that the buyers’ prior distributions are continuous and atomless, but not necessarily identical.
Theorem 2.
If each buyer’s prior distribution is regular, i.e., the virtual value $\phi(v)=v-\frac{1-F(v)}{f(v)}$ is increasing, then any Pareto point implementable by a randomized BIC auction is also implementable by a deterministic DSIC auction.
We prove the theorem in two steps. We first transform the allocation rule into a new one by only changing a single buyer’s interim allocation. We repeat this process $n$ times, and the mechanism with this new allocation rule yields weakly more welfare and revenue under regular distribution condition. In the second step, we show that the last allocation can be implemented in a deterministic DSIC mechanism. Since the original point is already Pareto point of the randomized BIC auction, the transformed auction implements the same point. The proof of Theorem 2 is deferred to Appendix A.2.
3.3 Equivalence of deterministic DSIC and randomized BIC auctions with respect to all (revenue,welfare) pairs for i.i.d distribution
In this subsection, we focus on the i.i.d case with a continuous and atomless distribution $F$ and utilize the quantile space. Since buyers’ private valuations are independently drawn from a common distribution $F$, we omit the subscripts on $v_{i}$ and $q_{i}$ and write them simply as $v$ and $q$. We provide sufficient conditions under which the range of (revenue, welfare) pairs achievable by deterministic auctions is equal to that of randomized BIC auctions.
Theorem 3.
If the prior distribution $F$ is third-order differentiable and satisfies the following condition, then any (revenue, welfare) pair that is implementable by a randomized auction is also implementable by a deterministic DSIC auction.
$v^{\prime\prime}(q)\geq 0,\quad\hat{\phi}^{\prime\prime}(q)\leq 0,\quad\forall q.$ (1)
It is not hard to verify that all uniform distributions and power-law distributions $F(v)=1-v^{-\alpha}$ with $\alpha\leq 1$, including the equal-revenue distributions, satisfy the condition.
The proof sketch is as follows: we first characterize the boundary of randomized BIC auctions, then we prove that each (revenue, welfare) pair on the boundary can be implemented by a deterministic auction (see Lemmas 1 - 2). Finally we prove that each (revenue, welfare) pair strictly between the revenue-maximizing boundary and the revenue-minimizing boundary at the same welfare level can be implemented by a deterministic auction (see Lemma 4). We begin by analyzing the boundary of (revenue, welfare) pairs for randomized BIC auctions. For convenience, we assume that $F$ has full support on $[0,1]$. By (maskin1984optimal), it is without loss of generality to only consider symmetric auctions. Define $P(q)=\int_{q}^{1}\hat{x}(t)\mathrm{d}t$ and we have
$\textrm{WEL}=n\int_{0}^{1}v(q)\hat{x}(q)\mathrm{d}q=-n\int_{0}^{1}v(q)\mathrm{d}P(q)=n\int_{0}^{1}v^{\prime}(q)P(q)\mathrm{d}q,$
$\textrm{REV}=n\int_{0}^{1}\hat{\phi}(q)\hat{x}(q)\mathrm{d}q=-n\int_{0}^{1}\hat{\phi}(q)\mathrm{d}P(q)=n\hat{\phi}(0)P(0)+n\int_{0}^{1}\hat{\phi}^{\prime}(q)P(q)\mathrm{d}q.$
Note that $P^{\prime}(q)=-\hat{x}(q)\leq 0$ and $P^{\prime\prime}(q)=-\hat{x}^{\prime}(q)\leq 0$ for a randomized BIC auction. By the feasibility from Border’s theorem, $P(q)$ is bounded by $P(q)\leq\frac{1-q^{n}}{n}$. We fix $\textrm{WEL}=c$ where $c$ is a constant, and then compute the range of REV. Thus, the optimization problem is
$\max/\min$ $~~~~n\hat{\phi}(0)P(0)+n\int_{0}^{1}\hat{\phi}^{\prime}(q)P(q)\mathrm{d}q$ (2)
$s.t.$ $~~~~n\int_{0}^{1}v^{\prime}(q)P(q)\mathrm{d}q=c,$
$~~~~P^{\prime}(q)\leq 0,~P^{\prime\prime}(q)\leq 0,~P(q)\leq\frac{1-q^{n}}{n},~P(1)=0.$
We define a class of auctions called “piecewise auctions”. We show that any point on the boundary of the feasible range can be implemented by a piecewise auction.
Definition 1 (Piecewise auctions).
A (symmetric, randomized) BIC auction is a piecewise auction if it satisfies one of the following two conditions:
1. There exists $0\leq r_{1}<r_{2}\leq 1$ such that
  • $P^{\prime}(q)=0$, for $0<q\leq r_{1}$,
  • $P^{\prime}(q)=-k$, for $r_{1}<q\leq r_{2}$, where $k$ is a constant and $0\leq k\leq\frac{r_{2}^{n}-r_{1}^{n}}{n(r_{2}-r_{1})}$,
  • $P(q)=\frac{1-q^{n}}{n}$, for $r_{2}<q\leq 1$.
2. There exists $0\leq r_{1}<r_{2}\leq 1$ such that
  • $P^{\prime}(q)=0$, for $0<q\leq r_{1}$,
  • $P(q)=\frac{1-q^{n}}{n}$, for $r_{1}<q\leq r_{2}$,
  • $P^{\prime}(q)=\frac{1-r_{2}^{n}}{n(r_{2}-1)}$, for $r_{2}<q\leq 1$.
According to the piecewise auctions we defined, we introduce the following Lemmas. Lemma 1 shows that any maximum and any minimum solutions of optimization problem (2) must be a piecewise auction, so all boundary points of the $(\mathrm{REV},\mathrm{WEL})$ region induced by randomized BIC auctions are attained by piecewise auctions. Lemma 3 then constructs, for each piecewise auction, a deterministic DSIC mechanism whose total interim allocation at every valuation coincides with that of the piecewise auction; together with Lemma 2, which states that identical total interim allocations imply identical revenue and welfare, this implies that every boundary point can be implemented by a deterministic DSIC auction.
Lemma 1.
If the condition (1) is satisfied, then any maximum and any minimum solutions of optimization problem (2) are a piecewise auction.
Lemma 2.
For any common prior distribution, if two auctions with interim allocation profile $(x_{1}(v),x_{2}(v),\cdots,x_{n}(v))$ and $(y_{1}(v),y_{2}(v),\cdots,y_{n}(v))$ respectively satisfy for any $v$,
$\sum_{i=1}^{n}x_{i}(v)=\sum_{i=1}^{n}y_{i}(v),$
then the two auctions’ revenue and welfare are the same.
By Lemma 2, we can get Lemma 3.
Lemma 3.
All (revenue, welfare) pairs of piecewise auctions can be implemented by deterministic DSIC auctions.
Lemma 4 further shows that, for any fixed welfare level, piecewise auctions span a continuous interval in the revenue dimension, so all interior points on the same welfare line are also implemented by piecewise auctions (and hence by deterministic DSIC mechanisms). Therefore, the entire $(\mathrm{REV},\mathrm{WEL})$ region attainable by randomized auctions can be implemented by deterministic DSIC auctions, which completes the proof of Theorem 3.
Lemma 4.
One piecewise auction can continuously transfer to another through a sequence of piecewise auctions with the welfare fixed.
4 Implementations with respect to interim allocations
4.1 Two buyers
In this subsection, we analyze the necessary and sufficient condition such that an interim allocation profile can be implemented by a deterministic DSIC auction, also known as Border’s Theorem in the deterministic case. We focus on the two-buyer case, with continuous and atomless prior distributions, which are not necessarily symmetric. The analysis is conducted in the quantile space.
Theorem 4.
An strictly increasing interim allocation profile $(\hat{x}_{1}(q),\hat{x}_{2}(q))$ is implementable by a deterministic DSIC auction if and only if
$\forall q,~\hat{x}_{2}(\hat{x}_{1}(q))\leq q.$ (3)
Moreover, if the item is always sold for the deterministic DSIC auction, then (3) turns into
$\forall q,~\hat{x}_{2}(\hat{x}_{1}(q))=q.$ (4)
The proof sketch is as follows: we first construct a corresponding three-coloring $C$ of the quantile plane, where $C\in\{0,1,2\}$, indicating that the item is allocated to no one, buyer 1 or buyer 2. Then, we rearrange a new coloring $C^{\prime}$, ensuring that the new interim allocations for both buyer 1 and buyer 2 remain unchanged. The resulting interim allocation profile can then be implemented by a deterministic DSIC auction. Furthermore, the quantile plane is partitioned by two monotone curves: $q_{2}=\hat{x}_{1}(q_{1})$ and $q_{1}=\hat{x}_{2}(q_{2})$. Finally, we establish that $\forall q,\hat{x}_{2}(\hat{x}_{1}(q))\leq q$.
Nonequivalence of deterministic BIC and DSIC auctions with respect to interim allocations
Based on Theorem 4, we can prove the nonequivalence between deterministic BIC and DSIC auctions with respect to interim allocations. A straightforward approach is to consider the sum of interim allocations given a valuation $v$.
Corollary 1.
There exists an interim allocation profile $(x_{1}(v),\cdots,x_{n}(v))$ that is implementable by a deterministic BIC auction, but there is no deterministic DSIC auction with interim allocation profile $(y_{1}(v),\cdots,y_{n}(v))$ such that for any $v$,
$\sum_{i=1}^{n}x_{i}(v)=\sum_{i=1}^{n}y_{i}(v).$
Next, we outline the idea of the proof. We consider the two-buyer case and analyze the following randomized BIC auction:
  • with probability $p$, run VCG auction.
  • with probability 1p1-p, send the item for free to buyer 1.
We know that the item is always sold. By Theorem 3, we have $\hat{x}_{2}(\hat{x}_{1}(q))=q$ for all $q$. This allows us to derive a series of interim allocations under a deterministic DSIC auction. However, the interim allocation may exceed 1, thereby proving the corollary.
4.2 Three buyers with constant interim allocation
Theorem 4 becomes extremely complex when extended to cases with three or more buyers. Suppose there are three buyers with constant interim allocations, i.e. $x_{i}(v)=c_{i},\forall v,i=1,2,3$. We provide a necessary and sufficient condition for interim allocations to be implemented by a deterministic DSIC auction.
Theorem 5.
The constant interim allocation profile for 3 buyers is implementable by a deterministic DSIC auction if and only if
$1-c_{1}-c_{2}-c_{3}\geq 2\sqrt{c_{1}c_{2}c_{3}}.$ (5)
The key insight stems from how deterministic DSIC constraints shape the allocation space. Since deterministic DSIC requires each buyer’s allocation to depend on critical values, we can model allocations through non-overlapping regions in the valuation space. For three buyers, we partition the three-dimensional space into regions A, B, and C where buyers 1, 2, and 3 win respectively. The non-overlapping nature of these regions creates geometric constraints on their projections. Thus, we obtain a series of inequalities and derive the necessity. For sufficiency, when the condition holds, we explicitly construct deterministic allocation rules. This construction ensures DSIC while satisfying the interim allocation probabilities.
5 Discussion
In this paper, we investigated the implementation power of deterministic auctions in single-item settings, focusing on two key aspects: (revenue, welfare) pairs and interim allocation rules. Our results highlight the distinctions between deterministic BIC and deterministic DSIC auctions. Several promising directions for future research emerge from this work. First, while we identified specific distributional conditions under which deterministic DSIC auctions achieve the same (revenue, welfare) pairs as randomized auctions, it would be valuable to explore whether these conditions can be relaxed or if other classes of distributions exhibit similar properties. Second, our characterization of interim allocation rules for deterministic DSIC auctions currently applies only to two and three buyers. Generalizing this to more buyers, even under simplifying assumptions (e.g., constant interim allocations), would be a significant step forward. By addressing these open questions, future research can further illuminate the boundaries of deterministic mechanism design and its applicability across diverse economic and computational settings.
Appendix A OMITTED PROOFS FROM SECTION 3
A.1 Experimental Results Regarding Pareto Points
Only partial figure is provided here since the complete figure is too large.
Figure 1: Pareto points in the area ($\textrm{WEL}>96$ and $\textrm{REV}>80$). Each blue point represents a Pareto point. The red line characterizes the boundary. The horizontal axis represents welfare, and the vertical axis represents revenue.
A.2 Proof of Theorem 2
Proof.
Fix a randomized auction with interim allocation profile $\{\hat{x}_{i}(q_{i})\}_{i=1}^{n}$ in the quantile space. We first transform the allocation rule into a new one by only changing a single buyer’s interim allocation. In a single process, we keep all buyer’s interim allocation rule except $i$ and increase buyer $i$’s interim allocation. For any quantile profile $\mathbf{q}=(q_{i},\mathbf{q}_{-i})\in[0,1]^{n}$, define function
$h_{i}(q_{i},\mathbf{q}_{-i})=1-\prod_{j}q_{j}-\sum_{j\neq i}\int^{1}_{q_{j}}\hat{x}_{j}(r_{j})\mathrm{d}r_{j}.$
The first negative term $\prod_{j}q_{j}$ is linear in $q_{i}$, and the second negative term $\sum_{j\neq i}\int^{1}_{q_{j}}\hat{x}_{j}(r_{j})\mathrm{d}r_{j}$ is independent of $q_{i}$. Thus, for any fixed $\mathbf{q}_{-i}$, the function $h_{i}(q_{i},\mathbf{q}_{-i})$ is linear in $q_{i}$. Now define
$\underline{h}_{i}(q_{i})=\min_{\mathbf{q}_{-i}}h_{i}(q_{i},\mathbf{q}_{-i}).$
Since $\underline{h}_{i}(q_{i})$ is the pointwise minimum of linear functions of $q_{i}$, it is concave in $q_{i}$. It is also decreasing and satisfies
$\underline{h}_{i}(1)=0,\qquad\underline{h}_{i}(0)=1-\sum_{j\neq i}\int_{0}^{1}\hat{x}_{j}(r_{j})\mathrm{d}r_{j}.$
The total interim allocation of buyer $i$ is $\int_{0}^{1}\hat{x}_{i}(r_{i})\mathrm{d}r_{i}$ lies in the interval $\underline{h}_{i}(1),\underline{h}_{i}(0)]$. Hence there exists a threshold $s_{i}^{*}\in[0,1]$ such that $\int_{0}^{1}\hat{x}_{i}(r_{i})\mathrm{d}r_{i}=\underline{h}_{i}(s_{i}^{*})$. Using this $s_{i}^{*}$, we define a new interim allocation for buyer $i$:
$\hat{x}_{i}^{*}(r_{i})$ $=\begin{cases}-\underline{h}_{i}^{\prime}(r_{i}),\quad&r_{i}\in[s_{i}^{*},1];\\ 0,\quad&r_{i}\in[0,s_{i}^{*}),\end{cases}$
where $\underline{h}_{i}^{\prime}$ is the (almost everywhere defined) derivative of the concave function $\underline{h}_{i}$. Because $\underline{h}_{i}$ is concave, $\underline{h}_{i}^{\prime}$ is nonincreasing, so $\underline{h}_{i}^{\prime}$ is nondecreasing. Since $\underline{h}_{i}$ is a concave envelope of a function with slopes in [1,0][-1,0], its derivative (where it exists) must also lie in [1,0][-1,0], and hence $\underline{h}^{\prime}_{i}\in[0,1]$. Therefore $\hat{x}_{i}^{*}(r_{i})$ is monotone in $r_{i}$. Moreover,
$\int_{0}^{1}\hat{x}_{i}^{*}(r_{i})\mathrm{d}r_{i}=\int_{s_{i}^{*}}^{1}-\underline{h}_{i}^{\prime}(r_{i})\mathrm{d}r_{i}=\underline{h}_{i}(s_{i}^{*})-\underline{h}_{i}(1)=\underline{h}_{i}(s_{i}^{*})=\int_{0}^{1}\hat{x}_{i}(r_{i})\mathrm{d}r_{i},$
so buyer $i$’s total interim allocation is preserved. By the definition of $\underline{h}_{i}$, for any threshold profile $(s_{i},\mathbf{s}_{-i})$, we have
$\int^{1}_{s_{i}}\hat{x}_{i}^{*}(r_{i})\mathrm{d}r_{i}+\sum_{j\neq i}\int^{1}_{s_{j}}\hat{x}_{j}(r_{j})\mathrm{d}r_{j}\leq 1-\prod_{j}s_{j},$
so the Border feasibility constraints continue to hold when we replace $\hat{x}_{i}$ by $\hat{x}_{i}^{*}$ and leave all other buyers’ interim allocations unchanged.
Next, we compare welfare and revenue from buyer $i$. Let $v_{i}(q_{i})$ denote the value of buyer $i$ at quantile $q_{i}$. The expected welfare contribution of buyer $i$ can be written as
$\int^{1}_{q_{i}=0}v_{i}(q_{i})\hat{x}_{i}(q_{i})\mathrm{d}q_{i}=\int^{1}_{q_{i}=0}\left(\int_{t=q_{i}}^{1}\hat{x}_{i}(t)\mathrm{d}t\right)\mathrm{d}v_{i}(q_{i}).$
Because $\hat{x}_{i}^{*}$ is obtained from a transformation that shifts buyer $i$’s allocation probability toward higher quantiles (and thus higher values) while preserving the total allocation mass, and because $v_{i}(q_{i})$ is increasing in $q_{i}$, we have
$\int^{1}_{q_{i}=0}\left(\int_{t=q_{i}}^{1}\hat{x}_{i}(t)\mathrm{d}t\right)\mathrm{d}v_{i}(q_{i})\leq\int^{1}_{q_{i}=0}\left(\int_{t=q_{i}}^{1}\hat{x}^{*}_{i}(t)\mathrm{d}t\right)\mathrm{d}v_{i}(q_{i})=\int^{1}_{q_{i}=0}v_{i}(q_{i})\hat{x}_{i}^{*}(q_{i})\mathrm{d}q_{i}.$
Thus buyer $i$’s contribution to welfare weakly increases.
For revenue, we use the Myersonian representation in terms of the virtual value $\phi_{i}$. Under regularity, $\phi_{i}$ is increasing in the value and hence increasing in the quantile. The same monotone reshaping that improves welfare also weakly increases the expected virtual surplus from buyer $i$, and therefore weakly increases revenue from buyer $i$.
We apply this transformation sequentially for buyers $i=1,2,\cdots,n$: first obtain $\hat{x}_{1}^{*}$ and $s_{1}^{*}$ from the original profile, then obtain $\hat{x}_{2}^{*}$ and $s_{2}^{*}$ using $\hat{x}_{1}^{*}$ and $\hat{x}_{j}$ for $j\geq 2$, and so on, until we obtain $\hat{x}_{n}^{*}$ and $s_{n}^{*}$ from $(\hat{x}_{1}^{*},\cdots,\hat{x}_{n-1}^{*},\hat{x}_{n})$. The resulting interim allocation profile $\hat{\mathbf{x}}^{*}=(\hat{x}_{1}^{*},\cdots,\hat{x}_{n}^{*})$ is Border-feasible, monotone for each buyer, and yields weakly higher welfare and revenue than the original randomized auction. This completes the first step of the proof.
The second step shows that $\hat{\mathbf{x}}^{*}$ can be implemented by a deterministic DSIC mechanism. We have the following lemma.
Lemma 5.
$\hat{\mathbf{x}}^{*}$ can be implemented in a deterministic DSIC mechanism.
Proof.
Fix a buyer $i$ and a quantile $q_{i}$. Under the new allocation rule $\hat{\mathbf{x}}^{*}$, Border’s constraint is still tight for some profiles of the other buyers’ quantiles $\mathbf{q}_{-i}$. First, we show that all such tight profiles for $(i,q_{i})$ are coordinatewise ordered. Suppose not, we assume there exist two tight profiles $\mathbf{q}_{-i}^{1}$ and $\mathbf{q}_{-i}^{2}$ such that for some buyers $j_{1}$ and $j_{2}$, we have $q^{1}_{j_{1}}<q^{2}_{j_{1}}\quad\text{and}\quad q^{1}_{j_{2}}>q^{2}_{j_{2}}$. Construct a new profile that mixes these two coordinates, that is $q_{j_{1}}=\frac{q^{1}_{j_{1}}+q^{2}_{j_{1}}}{2}$ and $q_{j_{2}}=\frac{q^{1}_{j_{2}}+q^{2}_{j_{2}}}{2}$, keeping the others fixed. Because the Border constraint is tight at $(q_{i},\mathbf{q}_{-i}^{1})$, feasibility of $\hat{\mathbf{x}}^{*}$ forces the item to be allocated to buyer $j_{1}$ at this mixed profile. Because the Border constraint is also tight at $(q_{i},\mathbf{q}_{-i}^{2})$, it simultaneously forces the item to be allocated to buyer $j_{2}$. This contradicts single-item feasibility. Hence, for fixed $(i,q_{i})$, any two tight profiles $(\mathbf{q}_{-i}^{1},\mathbf{q}_{-i}^{2})$ must be coordinatewise comparable: either $\mathbf{q}_{-i}^{1}\leq\mathbf{q}_{-i}^{2}$ or $\mathbf{q}_{-i}^{2}\leq\mathbf{q}_{-i}^{1}$. Now, for each pair of buyers $i\neq j$ and each $q_{i}\in[s_{i}^{*},1]$, define
$\lambda_{i,j}(q_{i})=\max\{q_{j}\mid(q_{i},\mathbf{q}_{-i})\text{ makes Border's constraint tight.}\},q_{i}\in[s_{i}^{*},1].$
By the ordering above, this maximum is well-defined, and the profile $(q_{i},\lambda_{i,1}(q_{i}),\lambda_{i,2}(q_{i}),\cdots)$ is itself a tight quantile profile. Writing $\lambda_{i}(q_{i})=(\lambda_{i,j}(q_{i}))_{j\neq i}$, these tight profiles induce a set $\mathcal{L}$ of relations $l=(q_{1},\cdots,q_{n})$ among buyers’ types. Using the same “no-crossing” argument as before, one shows that the relations in $\mathcal{L}$ are themselves mutually ordered. For each relation $l=(q_{1},\cdots,q_{n})\in\mathcal{L}$, define its total score $c(l)=\sum_{i}q_{i}$. Then, for each buyer $i$ and each $q_{i}\in[s_{i}^{*},1]$, define the individual score
$c_{i}(q_{i})=\max\{c(l)\mid l=(q_{1},\cdots,q_{n})\in\mathcal{L}\},q_{i}\in[s_{i}^{*},1].$
When $q_{i}\in[0,s_{i}^{*})$, let $c_{i}(q_{i})=0$. Thus, each type $q_{i}$ is assigned a scalar priority $c_{i}(q_{i})$. We now specify a deterministic hierarchical allocation rule in quantile space. Given any profile $\mathbf{q}=(q_{1},\cdots,q_{n})$, we compute scores $c_{i}(q_{i})$ for all buyers. Next,
  • If $\max_{i}c_{i}(q_{i})=0$, the seller keeps the item.
  • If there is a unique non-zero highest score, the buyer with the highest score wins.
  • If there is multiple non-zero highest score, the highest buyer with the smallest index wins.
This rule is deterministic by construction. Moreover, since the scores $c_{i}(q_{i})$ are derived from ordered tight profiles, they are non-decreasing in $q_{i}$, and the hierarchical selection is monotone in each buyer’s own quantile. By the standard single-parameter characterization, this monotone deterministic allocation can be supported by appropriate critical-value payments, so it is DSIC. Finally, because the scores and relations are built exactly from the tight Border profiles of $\hat{\mathbf{x}}^{*}$, the interim allocation of this hierarchical mechanism coincides with $\hat{\mathbf{x}}^{*}$. Thus $\hat{\mathbf{x}}^{*}$ is implementable by a deterministic DSIC mechanism. ∎
A.3 Proof of Lemma 1
Proof.
We show that the solutions of optimization problem (2) actually take the forms depicted in Figure 2, which represents special cases of piecewise auctions.
(a) Maximum solution
(b) Maximum solution
(c) Minimum solution
Figure 2: Optimal Solutions of Optimization Problem (2)
First, we prove that in the maximum solution of optimization problem (2), the function $P(q)$ has the form in Figure 2(a) or 2(b), that is, $P(q)$ has the minimum stochastic order (for any feasible function $\hat{P}(q)$, $\int_{t}^{1}P(q)\mathrm{d}q\leq\int_{t}^{1}\hat{P}(q)\mathrm{d}q$, $\forall t\in[0,1]$).
We fix the value of $P(0)$ and suppose that there exists another feasible solution, denoted by $\hat{P}(q)$, which is different from $P(q)$. By concavity, we know that $\hat{P}(q)$ crosses $P(q)$ at exactly one point $Q$, occurring at $q=q_{0}$, as shown in Figure 3(a). Note that it is possible that the curve $P(q)$ and $\hat{P}(q)$ overlap.
(a) Maximum solution
(b) Minimum solution
Figure 3: Comparison Between $P(q)$ and $\hat{P}(q)$
Since $\int_{0}^{1}v^{\prime}(q)P(q)\mathrm{d}q=\int_{0}^{1}v^{\prime}(q)\hat{P}(q)\mathrm{d}q=\frac{c}{n}$, we have
$\int_{0}^{q_{0}}v^{\prime}(q)(P(q)-\hat{P}(q))\mathrm{d}q=\int_{q_{0}}^{1}v^{\prime}(q)(\hat{P}(q)-P(q))\mathrm{d}q.$ (6)
We know that $P(q)\geq\hat{P}(q),\forall 0\leq q\leq q_{0}$ and $P(q)\leq\hat{P}(q),\forall q_{0}\leq q\leq 1$. Together with $v^{\prime\prime}(q)\geq 0$, we have
$\int_{0}^{q_{0}}v^{\prime}(q)(P(q)-\hat{P}(q))\mathrm{d}q\leq\int_{0}^{q_{0}}v^{\prime}(q_{0})(P(q)-\hat{P}(q))\mathrm{d}q,$ (7)
$\int_{q_{0}}^{1}v^{\prime}(q)(\hat{P}(q)-P(q))\mathrm{d}q\geq\int_{q_{0}}^{1}v^{\prime}(q_{0})(\hat{P}(q)-P(q))\mathrm{d}q.$ (8)
Put equality (6), and inequalities (7) and (8) together, we have
$\int_{0}^{q_{0}}(P(q)-\hat{P}(q))\mathrm{d}q\geq\int_{q_{0}}^{1}(\hat{P}(q)-P(q))\mathrm{d}q.$
Similarly, due to $\hat{\phi}^{\prime\prime}(q)\leq 0$, the difference of revenue between the solution with $P(q)$ and $\hat{P}(q)$ is
$n\int_{0}^{q_{0}}\hat{\phi}^{\prime}(q)(P(q)-\hat{P}(q))\mathrm{d}q-n\int_{q_{0}}^{1}\hat{\phi}^{\prime}(q)(\hat{P}(q)-P(q))\mathrm{d}q$
$\geq$~ $n\hat{\phi}^{\prime}(q_{0})\cdot\left(\int_{0}^{q_{0}}(P(q)-\hat{P}(q))\mathrm{d}q-\int_{q_{0}}^{1}(\hat{P}(q)-P(q))\mathrm{d}q\right)$
$\geq$~ $0.$
Finally, we prove that $\hat{\phi}^{\prime}(q_{0})\geq 0$. If $\hat{\phi}^{\prime}(q_{0})<0$, then $\hat{\phi}^{\prime}(q)<0$ for all $q>q_{0}$ according to $\hat{\phi}^{\prime\prime}(q)\leq 0$. Thus, $\hat{\phi}(q_{0})>\hat{\phi}(1)=1$, which leads to a contradiction. So, we have $\hat{\phi}^{\prime}(q_{0})\geq 0$. Therefore, we solve the maximum solution.
For the minimum solution, we prove that the function $P(q)$ has the form in Figure 2(c), that is $P(q)$ has the maximum stochastic order. The proof is similar and omitted. ∎
A.4 Proof of Lemma 3
Proof.
For piecewise auctions satisfying the first conditions in Definition 1, the corresponding interim allocations are as follows:
  • $\hat{x}(q)=0$, for $0<q\leq r_{1}$,
  • $\hat{x}(q)=k$, for $r_{1}<q\leq r_{2}$,
  • $\hat{x}(q)=q^{n-1}$, for $r_{2}<q\leq 1$.
Define $A=\{i|q_{i}>r_{2}\}$, $B=\{i|r_{1}\leq q_{i}\leq r_{2}\}$ and construct the deterministic allocation rule as follows:
  • If $|A|\neq 0$, set $\hat{x}_{j}(q_{1},q_{2},\cdots,q_{n})=1$, where $j$ is the minimum element in $A$ such that $q_{j}=\max_{i\in A}\{q_{i}\}$, and set $\hat{x}_{i}(q_{1},q_{2},\cdots,q_{n})=0,\forall i\neq j$.
  • If $|A|=0$, $|B|>0$ and $\frac{r_{2}^{n-1}}{n}\leq k\leq\frac{r_{2}^{n}-r_{1}^{n}}{n(r_{2}-r_{1})}$:
    • If $1\in B$, then set $\hat{x}_{1}(q_{1},q_{2},\cdots,q_{n})=1$, set $\hat{x}_{i}(q_{1},q_{2},\cdots,q_{n})=0,\forall i>1$.
    • If $1\notin B$, suppose that $j$ is the minimum element in $B$,
        If $q_{1}\leq\lambda r_{1}$, set $\hat{x}_{j}(q_{1},q_{2},\cdots,q_{n})=1,\hat{x}_{i}(q_{1},q_{2},\cdots,q_{n})=0,\forall i\neq j$.
        If $\lambda r_{1}<q_{1}<r_{1}$, set $\hat{x}_{i}(q_{1},q_{2},\cdots,q_{n})=0,\forall i$. ($\lambda$ ranges from 0 to $1$.)
  • If $|A|=0$, $|B|>0$ and $k<\frac{r_{2}^{n-1}}{n}$:
    • If $1\in B$,
        If $q_{2}\leq\lambda r_{2}$, set $\hat{x}_{1}(q_{1},q_{2},\cdots,q_{n})=1,\hat{x}_{i}(q_{1},q_{2},\cdots,q_{n})=0,\forall i>1$.
        If $\lambda r_{2}<q_{2}\leq r_{2}$, set $\hat{x}_{i}(q_{1},q_{2},\cdots,q_{n})=0,\forall i$. ($\lambda$ ranges from 0 to $1$.)
    • If $1\notin B$, set $\hat{x}_{i}(q_{1},q_{2},\cdots,q_{n})=0,\forall i$.
  • If $|A|=0$ and $|B|=0$, set $\hat{x}_{i}(q_{1},q_{2},\cdots,q_{n})=0,\forall i$.
The payment rule is according to the payment identity.
It is obvious that $\hat{x}_{i}(q)=0,q<r_{1}$ and $\hat{x}_{i}(q)=q^{n-1},q>r_{2}$. We then analyze the case $r_{1}\leq q\leq r_{2}$.
If $\frac{r_{2}^{n-1}}{n}\leq k\leq\frac{r_{2}^{n}-r_{1}^{n}}{n(r_{2}-r_{1})}$, we have $\hat{x}_{1}(q)=Pr[q_{j}\leq r_{2},\forall j\neq 1]=r_{2}^{n-1}$. And $\forall i>1$,
$\hat{x}_{i}(q)=Pr[q_{j}<r_{1},\forall j<i]\cdot Pr[q_{j}<r_{2},\forall j>i]\cdot Pr[q_{1}<\lambda r_{1}|q_{1}<r_{1}]=\lambda r_{1}^{i-1}r_{2}^{n-i}.$
Thus, we have
$\sum_{i=1}^{n}\hat{x}_{i}(q)=r_{2}^{n-1}+\sum_{i=2}^{n}\lambda r_{1}^{i-1}r_{2}^{n-i}\in[r_{2}^{n-1},\frac{r_{2}^{n}-r_{1}^{n}}{r_{2}-r_{1}}].$
Therefore, there exists $\lambda\in[0,1]$ such that $\sum_{i=1}^{n}\hat{x}_{i}(q)=nk$.
If $k<r_{2}^{n-1}$, $\hat{x}_{i}(q)=0,\forall i>1$. And we have
$\hat{x}_{1}(q)=Pr[q_{j}\leq r_{2},\forall j\neq i]\cdot Pr[q_{2}\leq\lambda r_{2}|q_{2}\leq r_{2}]=\lambda r_{2}^{n-1}\in[0,r_{2}^{n-1}].$
Thus, there exists $\lambda\in[0,1]$ such that $\sum_{i=1}^{n}\hat{x}_{i}(q)=nk$.
For piecewise auctions satisfying the second conditions in Definition 1, the corresponding interim allocations are as follows:
  • $\hat{x}(q)=0$, for $0<q\leq r_{1}$,
  • $\hat{x}(q)=q^{n-1}$, for $r_{1}<q\leq r_{2}$,
  • $\hat{x}(q)=\frac{1-r_{2}^{n}}{n(1-r_{2})}$, for $r_{2}<q\leq 1$.
Define $A=\{i|q_{i}>r_{2}\}$, $B=\{i|r_{1}\leq q_{i}\leq r_{2}\}$ and construct the deterministic allocation rule as follows:
  • If $|A|>0$, set $\hat{x}_{j}(q_{1},q_{2},\cdots,q_{n})=1$ where $j$ is the minimum element in $A$, and set $\hat{x}_{i}(q_{1},q_{2},\cdots,q_{n})=0,\forall i\neq j$.
  • If $|A|=0$ and $|B|>0$, set $\hat{x}_{j}(q_{1},q_{2},\cdots,q_{n})=1$ where $j$ is the minimum element in $B$ such that $q_{j}=\max_{i\in B}\{q_{i}\}$, and set $\hat{x}_{i}(q_{1},q_{2},\cdots,q_{n})=0,\forall i\neq j$.
  • If $|A|=0$ and $|B|=0$, set $\hat{x}_{i}(q_{1},q_{2},\cdots,q_{n})=0,\forall i$.
The payment rule is according to the payment identity.
For all $q>r_{2}$, we have $\hat{x}_{i}(q)=Pr[q_{j}<r_{2},\forall j<i]=r_{2}^{i-1}$, thus $\sum_{i=1}^{n}\hat{x}_{i}(q)=\sum_{i=1}^{n}r_{2}^{i-1}=\frac{1-r_{2}^{n}}{1-r_{2}}$.
For all $r_{1}\leq q\leq r_{2}$, $\hat{x}_{i}(q)=Pr[q_{j}<q,\forall j\neq i]=q^{n-1}$ , thus $\sum_{i=1}^{n}\hat{x}_{i}(q)=nq^{n-1}$.
For all $q<r_{1}$, $\hat{x}_{i}(q)=0$.
So we have $\sum_{i=1}^{n}\hat{x}_{i}(q)=n\hat{x}(q)$. By Lemma 2, the (revenue, welfare) pairs of piecewise auctions is implementable by deterministic DSIC auctions. ∎
A.5 Proof of Lemma 2
Proof.
The revenue and the welfare for the auction with interim allocation profile $(x_{1}(v),\cdots,x_{n}(v))$ are
$\textrm{WEL}=\sum_{i=1}^{n}\int_{0}^{1}vx_{i}(v)f(v)\mathrm{d}v=\int_{0}^{1}vf(v)\sum_{i=1}^{n}x_{i}(v)\mathrm{d}v,$
$\textrm{REV}=\sum_{i=1}^{n}\int_{0}^{1}\phi(v)x_{i}(v)f(v)\mathrm{d}v=\int_{0}^{1}\phi(v)f(v)\sum_{i=1}^{n}x_{i}(v)\mathrm{d}v.$
Because welfare and revenue are only related to the interim allocation profile, $\sum_{i=1}^{n}x_{i}(v)=\sum_{i=1}^{n}y_{i}(v)$, thus two auctions’ revenue and welfare are the same. ∎
A.6 Proof of Lemma 4
Proof.
Fix a welfare level $c$. Let $\mathcal{M}^{a}$ and $\mathcal{M}^{b}$ be two piecewise auctions (Definition 1) with $\mathrm{WEL}(\mathcal{M}^{a})=\mathrm{WEL}(\mathcal{M}^{b})=c$. We construct a continuous transfer (via a sequence of piecewise auctions, all having welfare $c$) from each auction to a common canonical piecewise auction. Throughout, write the interim allocation as $\hat{x}(q)$, $q\in[0,1]$, so
$\mathrm{WEL}(\hat{x})=n\int_{0}^{1}v(q)\hat{x}(q)\mathrm{d}q,$
where $v(q)=F^{-1}(q)$.
Define the “reserve-type” interim allocation
$\hat{x}^{\rho}(q):=\mathbb{I}\{q>\rho\}q^{n-1}.$
Its welfare is
$W(\rho):=n\int_{\rho}^{1}v(q)q^{n-1}\mathrm{d}q,$
which is continuous and (weakly) decreasing in $\rho$. Hence there exists $\rho^{\star}\in[0,1]$ such that $W(\rho^{\star})=c$. Let $\mathcal{M}^{\star}$ denote the piecewise auction implementing $\hat{x}^{\rho^{\star}}$. Note $\hat{x}^{\rho^{\star}}$ is exactly a Condition 2 piecewise auction in Definition 1 with parameters $(r_{1},r_{2})=(\rho^{\star},1)$. It suffices to show: any piecewise auction $\mathcal{M}$ with welfare $c$ can be continuously transferred to $\mathcal{M}^{\star}$ through piecewise auctions of welfare $c$.
First, we discuss Condition 2 in Definition 1. A Condition 2 piecewise auction is characterized by parameters $0\leq r_{1}<r_{2}\leq 1$ and interim allocation
$\hat{x}(q)=\begin{cases}0,&q\leq r_{1};\\ q^{n-1},&r_{1}<q\leq r_{2};\\ \kappa(r_{2}):=\dfrac{1-r_{2}^{n}}{n(1-r_{2})},&r_{2}<q\leq 1.\end{cases}$
Define $r_{2}(t):=(1-t)r_{2}+t\text{ for }t\in[0,1]$. For each $t$, choose $r_{1}(t)\in[0,r_{2}(t))$ so that the resulting Condition 2 piecewise auction $\mathcal{M}_{t}$ satisfies $\mathrm{WEL}(\mathcal{M}_{t})=c$. This is well-defined and continuous in $t$ because, for fixed $r_{2}(t)$, welfare is a continuous and strictly decreasing function of $r_{1}$ (increasing $r_{1}$ removes positive allocation $q^{n-1}$ on $(r_{1},r_{2}(t)])$. Thus $t\mapsto r_{1}(t)$ is continuous, $\mathcal{M}_{t}$ stays within Condition 2, and $\mathcal{M}_{1}$ has $(r_{1}(1),r_{2}(1))=(\rho^{\star},1)$, i.e. $\mathcal{M}_{1}=\mathcal{M}^{\star}$.
Second, we discuss Condition 1 in Definition 1. A Condition 1 piecewise auction has parameters $0\leq r_{1}<r_{2}\leq 1$ and constant $k$ with
$0\leq k\leq k_{\max}(r_{1},r_{2}):=\frac{r_{2}^{n}-r_{1}^{n}}{n(r_{2}-r_{1})},$
and interim allocation
$\hat{x}(q)=\begin{cases}0,&q\leq r_{1},\\ k,&r_{1}<q\leq r_{2},\\ q^{n-1},&r_{2}<q\leq 1.\end{cases}$
Let $\mathcal{M}$ be such an auction with welfare $c$. Define
$T(r):=n\int_{r}^{1}v(q)q^{n-1}\mathrm{d}q,\quad(\text{tail welfare}).$
Since $\mathrm{WEL}(\mathcal{M})=T(r_{2})+nk\int_{r_{1}}^{r_{2}}v(q)\mathrm{d}q=c$ with $k\geq 0$, we have $T(r_{2})\leq c=T(\rho^{\star})$, hence $r_{2}\geq\rho^{\star}$. Now set $r_{2}(t):=(1-t)r_{2}+t\rho^{\star}$, so $r_{2}(t)\in[\rho^{\star},r_{2}]$. Keep $r_{1}$ fixed. For each $t$, define $k(t)$ as the unique value making welfare equal to $c$, we have
$k(t):=\frac{c-T(r_{2}(t))}{n\int_{r_{1}}^{r_{2}(t)}v(q)\mathrm{d}q}.$
This is well-defined because $T(r_{2}(t))\leq T(\rho^{\star})=c$ and $\int_{r_{1}}^{r_{2}(t)}v(q)\mathrm{d}q>0$. Moreover, feasibility $k(t)\leq k_{\max}(r_{1},r_{2}(t))$ holds because, for fixed $(r_{1},r_{2}(t))$, welfare is affine increasing in $k$ on the feasible interval $0,k_{\max}]$. Since $k(0)=k$ is feasible and $t\mapsto(r_{2}(t),k(t))$ is continuous, the equality $\mathrm{WEL}=c$ selects a feasible $k(t)\in[0,k_{\max}(r_{1},r_{2}(t))]$ for all $t$. Thus each $\mathcal{M}_{t}$ defined by $(r_{1},r_{2}(t),k(t))$ is a Condition 1 piecewise auction with welfare $c$, and $t\mapsto\mathcal{M}_{t}$ is continuous. At $t=1$, $r_{2}(1)=\rho^{\star}$ and $T(\rho^{\star})=c$, hence $k(1)=0$. Therefore
$\hat{x}_{M_{1}}(q)=\mathbb{I}\{q>\rho^{\star}\}q^{n-1}=\hat{x}^{\rho^{\star}},$
so $\mathcal{M}_{1}=\mathcal{M}^{\star}$.
Apply the above transfer process, we can transfer $\mathcal{M}^{a}$ continuously to $\mathcal{M}^{\star}$ through piecewise auctions with welfare $c$, and similarly transfer $\mathcal{M}^{b}$ to $\mathcal{M}^{\star}$. Concatenating the first transfer with the reverse of the second yields a sequence of piecewise auctions that continuously transfers $\mathcal{M}^{a}$ to $\mathcal{M}^{b}$ while keeping welfare fixed. ∎
A.7 The Proof of Theorem 3
Proof.
Fix a welfare level $c$. Consider the optimization problem 2 that maximizes (resp. minimizes) revenue over all feasible symmetric randomized auctions subject to $\mathrm{WEL}=c$. Let $R_{\max}(c)$ and $R_{\min}(c)$ denote the optimal values.
By Lemma 1, there exist optimal solutions to these two problems whose induced auctions are piecewise. Hence there are two piecewise auctions $\mathcal{M}^{\max}$ and $\mathcal{M}^{\min}$ such that
$\mathrm{WEL}(\mathcal{M}^{\max})$ $=\mathrm{WEL}(\mathcal{M}^{\min})=c,$
$\mathrm{REV}(\mathcal{M}^{\max})$ $=R_{\max}(c),\quad\mathrm{REV}(\mathcal{M}^{\min})=R_{\min}(c).$
By Lemma 3, each piecewise auction can be implemented by a deterministic DSIC mechanism with the same total interim allocation. By Lemma 2, matching total interim allocation implies matching expected welfare and expected revenue. Therefore, the two boundary points $(R_{\max}(c),c)$ and $(R_{\min}(c),c)$ are both achievable by deterministic DSIC auctions. Next, Lemma 4 implies that $\mathcal{M}^{\max}$ can be continuously transferred to $\mathcal{M}^{\min}$ through a sequence of piecewise auctions while keeping welfare fixed at $c$. Along this transfer, revenue varies continuously, so every intermediate revenue level $r\in[R_{\min}(c),R_{\max}(c)]$ is achieved by some piecewise auction $\mathcal{M}$ with $\mathrm{WEL}(\mathcal{M})=c$ and $\mathrm{REV}(\mathcal{M})=r$. Applying Lemmas 3 and 2 again, each such point $(r,c)$ is implementable by a deterministic DSIC auction. Finally, any randomized auction with welfare $c$ must have revenue in $R_{\min}(c),R_{\max}(c)]$ by definition of $R_{\min}(c)$ and $R_{\max}(c)$. Hence every $(\mathrm{REV},\mathrm{WEL})$ pair attainable by randomized auctions at welfare level $c$ is attainable by deterministic DSIC auctions. Since $c$ was arbitrary, the entire $(\mathrm{REV},\mathrm{WEL})$ region attainable by randomized auctions is attainable by deterministic DSIC auctions, proving Theorem 3. ∎
Appendix B OMITTED PROOFS FROM SECTION 4
B.1 Proof of Theorem 4
Proof.
Consider the plane
$A=\{(q_{1},q_{2}),0\leq q_{1}\leq 1,0\leq q_{2}\leq 1\}.$
We first prove the “only if” direction: for any deterministic DSIC auction with interim allocation profiles $(\hat{x}_{1}(q),\hat{x}_{2}(q))$, condition (3) holds. We construct a corresponding three-coloring $C$ of plane $A$: $C(q_{1},q_{2})\in\{0,1,2\}$, which means that the item is allocated to no one, buyer 1 or buyer 2.
By [nisan2007algorithmic], we know that if a DSIC auction is deterministic, then for any $i$, there exists a value $c_{i}(\mathbf{v}_{-i})$ that is independent of $v_{i}$, called the critical value, such that
  • If $v_{i}\geq c_{i}(\mathbf{v}_{-i})$, then $x_{i}(v_{i},\mathbf{v}_{-i})=1$, $p_{i}(v_{i},\mathbf{v}_{-i})=c_{i}(\mathbf{v}_{-i})$.
  • If $v_{i}<c_{i}(\mathbf{v}_{-i})$, then $x_{i}(v_{i},\mathbf{v}_{-i})=0$, $p_{i}(v_{i},\mathbf{v}_{-i})=0$.
Thus, for each quantile $q_{2}$ of buyer 2, there is a function $c_{1}(q_{2})$ which only depends on $q_{2}$, called the critical value for buyer 1, i.e, $\hat{x}_{1}(q,q_{2})=1$ if and only if $q\geq c_{1}(q_{2})$.
Next, we rearrange the coloring $C$ to $C^{\prime}$ with fixed interim allocations.
  • $C^{\prime}(q_{1},q_{2})=1$ for all $q_{2}\leq\hat{x}_{1}(q_{1})$,
  • $C^{\prime}(q_{1},q_{2})=0$ for all $q_{2}>\hat{x}_{1}(q_{1})$ and $C(q_{1},q_{2})=1$.
  • $C^{\prime}(q_{1},q_{2})=2$ for all $q_{2}>\hat{x}_{1}(q_{1})$ and $C(q_{1},q_{2})=2$.
Figure 4: Rearrangment on three-coloring of deterministic auctions
In fact, the rearrangement moves down all the area of buyer 1. Next, we show that the interim allocations keep unchanged. For buyer 1 with any quantile $q_{1}$, the new interim allocation is $\int_{q_{2}}\mathbf{I}[C^{\prime}(q_{1},q_{2})=1]\mathrm{d}q_{2}=\hat{x}_{1}(q_{1})$, where $\mathbf{I}$ is the indicator function. So the interim allocations remain the same for buyer 1.
For buyer 2, we prove that all areas that belong to buyer 2 remain the same. Otherwise, suppose that there is a point $(q_{1},q_{2})$ such that $C(q_{1},q_{2})=2$ and $C^{\prime}(q_{1},q_{2})\neq 2$. Then by the construction, it is easy to see that $q_{2}<\hat{x}_{1}(q_{1})$. By DSIC, we have $C(q_{1},q)=2$ for all $q>q_{2}$. Therefore, we have $\hat{x}_{1}(q_{1})\leq 1-\int_{q_{2}}^{1}\mathbf{I}[C(q_{1},q)=2]\mathrm{d}q_{2}=q_{2}<\hat{x}_{1}(q_{1})$, contradiction. (Here we ignore the case where $q_{2}=\hat{x}_{1}(q_{1})$ is the minimum value of $q_{2}$ such that $C(q_{1},q_{2})=2$, of which the probability is measured zero).
Do the same rearrangement for buyer 2, as shown in Figure 4, the area belongs to buyer 1 and buyer 2 are partitioned by two separate monotone curves, $q_{2}=\hat{x}_{1}(q_{1})$ and $q_{1}=\hat{x}_{2}(q_{2})$. Since the two curves does not cross each other, (it is possible that the two curves overlap. The results still hold as the overlapped area is measured zero), we have $\forall q,~\hat{x}_{2}(\hat{x}_{1}(q))\leq q$.
If the item is always sold, then no point in plane $A$ is colored 0. So the two curves $q_{2}=\hat{x}_{1}(q_{1})$ and $q_{1}=\hat{x}_{2}(q_{2})$ totally overlap (the colored 0 area does not exists). It follows that $\forall q,~\hat{x}_{2}(\hat{x}_{1}(q))=q$.
For the “if” direction, given (3), we can simply construct the DSIC auction by using the two curves, $q_{2}=\hat{x}_{1}(q_{1})$ and $q_{1}=\hat{x}_{2}(q_{2})$, to partition the plane $A$. The condition (3) guarantees that the two curves do not cross each other, thus the construction is feasible. Moreover, if (4) holds, then the two curves fully overlap, thus all points are colored either 1 or 2. ∎
B.2 Proof of Corollary 1
Proof.
Consider the 2-buyer case and the following randomized BIC auction:
  • with probability $p$, run VCG auction.
  • with probability 1p1-p, send the item for free to buyer 1.
We analyze it in the quantile space. Then, for any quantile $q=F(v)$, the sum of the interim allocation is 2pq+1p2pq+1-p. We assume that there exists a deterministic DSIC auction with interim allocation profile $(\hat{x}_{1}(q),\hat{x}_{2}(q))$ such that
$\hat{x}_{1}(q)+\hat{x}_{2}(q)=2pq+1-p.$
Since
$\int_{0}^{1}(\hat{x}_{1}(q)+\hat{x}_{2}(q))\mathrm{d}q=\int_{0}^{1}(2pq+1-p)\mathrm{d}q=1,$
so the item is always sold. By Theorem  3, we have $\hat{x}_{2}(\hat{x}_{1}(q))=q$ for all $q$. Since $\hat{x}_{1}(0)+\hat{x}_{2}(0)=1-p$, either $\hat{x}_{1}(0)$ or $\hat{x}_{2}(0)$ must be equal to 0. Otherwise, if $\hat{x}_{1}(0)>0$ and $\hat{x}_{2}(0)>0$, then $\hat{x}_{2}(\hat{x}_{1}(0))\geq\hat{x}_{2}(0)>0$, contradiction. Thus, we without loss of generality to assume that $\hat{x}_{1}(0)=1-p$ and $x_{2}(0)=0$.
Therefore, we can get a series of sequence:
$u_{n}+u_{n+2}=2pu_{n+1}+1-p,u_{1}=0,u_{2}=1-p,$
and $\hat{x}_{1}(u_{n})=u_{n+1}$, $\hat{x}_{2}(u_{n})=u_{n-1}$.
So if the sequence $\{u_{n}\}$ does not exact reach to 1, then the interim allocation is not feasible. It is not hard to check that when $p=0.75$, $u_{5}=1.0312>1$. The corollary is proved. ∎
B.3 Proof of Theorem 5
Proof.
We first prove the “only if” direction: given constant interim allocation rule $x_{1},x_{2},x_{3}$ in a deterministic DSIC auction, then condition (5) holds.
Due to the property of deterministic DSIC and and constant interim allocation rule, we have the following statement:
Except for a zero measure of $(b,c)\in[0,1]^{2}$, either $x_{1}(v,b,c)=0,\forall v$ or $x_{1}(v,b,c)=1,\forall v$.
It means that a buyer’s allocation rule does not depend on his own value. In the following proof we ignore the unsatisfactory set which is measured 0 and has no effect to results.
Define three valuation domain sets: $A=\{(b,c):x_{1}(*,b,c)=1\}$, $B=\{(a,c):x_{2}(a,*,c)=1\}$ and $C=\{(a,b):x_{3}(a,b,*)=1\}$. Each set represents two players’ valuation profile when the mechanism allocates the item to the remaining player.
Define the projection of set A in each dimension as $A_{b}=\{b:\exists c~s.t.~(b,c)\in A\}$ and $A_{c}=\{c:\exists b~s.t.~(b,c)\in A\}$. Define $B_{a},B_{c},C_{a},C_{b}$ similarly.
We next claim that $B_{a}\cap C_{a}=\emptyset$, $A_{b}\cap C_{b}=\emptyset$, $A_{c}\cap B_{c}=\emptyset$. Suppose $B_{a}\cap C_{a}\neq\emptyset$, let $a^{*}$ be an element in $B_{a}\cap C_{a}$. Assume $(a^{*},c^{*})\in B$ and $(a^{*},b^{*})\in C$. By definition of $B$, $x_{2}(a^{*},b^{*},c^{*})=1$, while by definition of $C$, $x_{3}(a^{*},b^{*},c^{*})=1$ which leads a contradiction. So $B_{a}$ and $C_{a}$ do not overlap. Same argument applies for $A_{b}$ and $C_{b}$, $A_{c}$ and $B_{c}$.
Define $y_{1}=|A_{b}|$, $y_{2}=|B_{c}|$, and $y_{3}=|C_{a}|$. Then we have $|C_{b}|\leq 1-y_{1}$, $|A_{c}|\leq 1-y_{2}$ and $|B_{a}|\leq 1-y_{3}$ (since $A_{b}\cap C_{b}=\emptyset$).
The interim allocation $c_{1}=|A|$ is bounded by $|A_{b}||A_{c}|\leq y_{1}(1-y_{2})$, since $A_{b}$ and $A_{c}$ are projections of $A$.
Thus, the interim allocation profile $c_{1},c_{2},c_{3}$ is implementable implies that there exists $y_{1},y_{2},y_{3}\in[0,1]$ such that
$c_{1}\leq y_{1}(1-y_{2}),$
$c_{2}\leq y_{2}(1-y_{3}),$
$c_{3}\leq y_{3}(1-y_{1}).$
We have
$c_{2}\leq y_{2}(1-y_{3})\leq(1-\frac{c_{1}}{y_{1}})(1-\frac{c_{3}}{1-y_{1}}).$
That is,
$(1-c_{2})y_{1}^{2}+(c_{2}-c_{1}+c_{3}-1)y_{1}+c_{1}(1-c_{3})\leq 0.$
Due to the formula above has a solution for $y_{1}$, we have
$(c_{2}-c_{1}+c_{3}-1)\geq 4c_{1}(1-c_{2})(1-c_{3}),$
That is,
$(c_{1}+c_{2}+c_{3}-1)^{2}\geq 4c_{1}c_{2}c_{3}.$
Note that
$1$ $\geq\mathbb{E}_{v_{1},v_{2},v_{3}}[x_{1}(v_{1},v_{2},v_{3})+x_{2}(v_{1},v_{2},v_{3})+x_{3}(v_{1},v_{2},v_{3})]$
$=\mathbb{E}_{v_{1}}[x_{1}(v_{1})]+\mathbb{E}_{v_{2}}[x_{2}(v_{2})]+\mathbb{E}_{v_{3}}[x_{3}(v_{3})]$
$=c_{1}+c_{2}+c_{3}.$
Consequently, we have
$1-c_{1}-c_{2}-c_{3}\geq 2\sqrt{c_{1}c_{2}c_{3}}.$
For the “if” direction: define $\Delta=(c_{1}+c_{2}+c_{3}-1)^{2}-4c_{1}c_{2}c_{3}$ and suppose $\Delta\geq 0$. We choose $y_{1}=\frac{1+c_{1}-c_{2}-c_{3}+\sqrt{\Delta}}{2(1-c_{2})}$, $y_{2}=\frac{1+c_{2}-c_{1}-c_{3}+\sqrt{\Delta}}{2(1-c_{3})}$, $y_{3}=\frac{1+c_{3}-c_{2}-c_{1}+\sqrt{\Delta}}{2(1-c_{1})}$.
Then, we construct an auction with the following allocation rule, which is deterministic DSIC.
  • $x_{1}(v_{1},v_{2},v_{3})=1$ if and only if $v_{2}\leq y_{1}$ and $v_{3}>y_{2}$.
  • $x_{2}(v_{1},v_{2},v_{3})=1$ if and only if $v_{3}\leq y_{2}$ and $v_{1}>y_{3}$.
  • $x_{3}(v_{1},v_{2},v_{3})=1$ if and only if $v_{1}\leq y_{3}$ and $v_{2}>y_{1}$.
Finally, we have to prove that “Except for a zero measure of $(b,c)\in[0,1]^{2}$, either $x_{1}(v,b,c)=0,\forall v$ or $x_{1}(v,b,c)=1,\forall v$”:
According to the definition of deterministic DSIC auction, given $(b,c)$, there exist a critical value $cr(b,c)$ such that
  • $x_{1}(v,b,c)=0$, if $v<cr(b,c)$,
  • $x_{1}(v,b,c)=1$, if $v>cr(b,c)$,
  • $x_{1}(v,b,c)=0$ or $x_{1}(v,b,c)=1$, if $v=cr(b,c)$.
Define $A_{1}=\{(b,c)|cr(b,c)=0\}$, $A_{2}=\{(b,c)|cr(b,c)<1\}$. Clearly $A_{1}\subseteq A_{2}$. Note that
$c_{1}=x_{1}(0)=E_{b,c}[x_{1}(0,b,c)]\leq|A_{1}|\leq|A_{2}|\leq E_{b,c}[x_{1}(1,b,c)]=x_{1}(1)=c_{1}.$
Hence $|A_{1}|=|A_{2}|$, and since $A_{1}\subseteq A_{2}$, we must have $|A_{2}\backslash A_{1}|=0$. But $A_{2}\backslash A_{1}=\{(b,c):0<cr(b,c)<1\}$. Therefore, except for a zero-measure set of $(b,c)$, we have $cr(b,c)\in\{0,1\}$, which implies $x_{1}(\cdot,b,c)$ is constant in $v$ (up to the boundary point $v=0$ or $v=1$, which does not affect interim allocations). This completes the proof. ∎